Find the % yield of cisplatin which is formed in the result of when 2 moles of ammonia react with potassium tetrachloro platinate
Determine the mass of nh3 that is needed to complete react with platinate and the amount of cisplatin that form?
An experiment begin with 20g of platinate
And 12.36 actual yield of cisplatin.
1
Expert's answer
2020-06-15T14:41:47-0400

n(K2PtCl4) = 20 / 415 = 0.0482 mol

n(Pt(NH3)2Cl2) = 12.36 / 300 = 0.0412 mol

"\alpha" = 0.0412 / 0.0482 × 100% = 85.5%

m2(NH3) = 0.0482 × 17 = 0.82 g

m2(Pt(NH3)2Cl2) = 0.0482 × 300 = 14.46 g